My problem asks me this:
Solve each system of linear equations.
y=x+2
y=2x
Do I use the xy chart to help work this problem and where do I go from there?
The problem has 2 variables and I don't know how to bring it down to one variable.
Please help me!!

Found 2 solutions by ankor@dixie-net.com, nabla:
Answer by ankor@dixie-net.com(22740)

Solve each system of linear equations.
y=x+2
y=2x
:
Look at the 2nd equation, that means any place in the problem that y occurs,
you can replace it with 2x, right?
:
Take the 1st equation, replace y with 2x, then solve for x
2x = x + 2
2x - x = +2, subtracted x from both sides
x = 2
:
You can use either equation to find y. Replace x with 2, using the 2nd equation:
y = 2(2)
y = 4
:
This is called the substitution method, there are other methods you can use
:
Check solution in 1st equation, by replacing x with 2 and y with 4, to see if equality reigns:
y = x + 2
4 = 2 + 2; confirms your solutions, did this help?

Answer by nabla(475)

You can actually just subtract the first equation from the second in this case.
This gives:
0=x-2
2=x

And if x=2, y=4 by back substitution.


You could also solve by substitution, graphing, or other techniques from linear algebra (such as Cramer's Rule after some slight rearrangement).